OpenStudy (anonymous):

A jar contains 26 coins totaling $2.55 in nickels and dimes. How many dimes are in the jar?

OpenStudy (anonymous):

Identify variables: n=nickels d=dimes Formulate two equations: n+d=26 0.1d+0.05n=2.55 Solve: 10d+5n=255 n=26-d Substitute: 10d+5(26-d)=255 10d+130-5d=255 5d=125 d=25 25 dimes are in the jar